# Proving the Limited Scalability of Centralized Distributed Optimization via a New Lower Bound Construction

### v1

@@ 2.3 Failure of the previous construction in the homogeneous setting
[…]
@@ 3 A New “Worst-Case” Function
In this section, we give a less technical description of our lower bound construction and the main theorem from Section B. Instead of (6), we propose to use another “worst-case” function. For any $T,K\in\mathbb{N},$ and $e\geq a>1,$ we define the function $F_{T,K,a}\,:\,\mathbb{R}^{T}\to\mathbb{R}$ such that
@@
[…]
@@
and $x_{0}=\dots=x_{-K+1}\equiv 1.$ The main modification is that instead of the block $-\Psi(x_{i-1})\Phi(x_{i}),$ we use $-\Psi_{a}(x_{i-K})\dots\Psi_{a}(x_{i-2})\Psi_{a}(x_{i-1})\Phi(x_{i})$ (ignore $a$ for now). In the previous approach, it was sufficient for a worker to have $x_{i-1}\neq 0$ to discover the next $i$th coordinate. However, in our new construction, the worker needs $x_{i-1}\neq 0,x_{i-2}\neq 0,\dots,x_{i-K+1}\neq 0$ for that. With this modification, it is not sufficient for the “luckiest” worker to get the non-zero $i-1$th coordinate to discover the next coordinate: to progress to the $i$th coordinate, the worker should also have non-zero $i-2$th, …, $i-K+1$th coordinates.
@@
    For all $x\in\mathbb{R}^{T},$ $\textnormal{prog}^{1}(\nabla F_{T,K,a}(x))\leq\max\{\textnormal{prog}^{K}(x)+1,\textnormal{prog}^{1}(x)\}.$
@@
Lemma 3.2 (Lemmas B.6, B.7, and B.8).
The function $F_{T}$ satisfies:
@@
Figure 1: The functions $\Psi_{a}(x)$ and $\Gamma(x)$, along with their derivatives $\Phi^{\prime}_{a}(x)$ and $\Gamma^{\prime}(x)$. The plots of $\Phi(x)$ and $\Phi^{\prime}(x)$ are shown in (Carmon et al., 2020).
@@ 4 Lower Bound with Server-to-Worker (S2W) Communication
[…]
@@
[…]
The main observation in Theorems 4.1 and 4.2 is that it is not possible to scale both $d$ and $\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ by more than $\log^{4}(n+1)$ and $\log^{6}(n+1),$ respectively. Asymptotically, this scaling is significantly worse than the linear $n$ and square-root $\sqrt{n}$ scalings discussed in Section 1.1. For instance, if $n=10{,}000$ and $d$ is increased by a factor of $10$, we have to increase $n$ by a factor of $10^{3}$ (two factors more) to ensure that $\nicefrac{{\tau_{\textnormal{s}}d}}{{\log^{4}(n+1)}}$ does not change.
4.1 Proof Sketch
@@
Next, we define two sets of random variables: (i) let $\eta_{1,i,k}$ denote the number of stochastic gradient computations until the first moment when a coordinate is not zeroed out in the stochastic gradient oracle (see (26)), after the moment when the $(k-1)$th coordinate is no longer zeroed out in worker $i$; (ii) let $\mu_{1,i,k}$ be the number of received coordinates until the moment when the last received coordinate belongs to $[K]$, after the $(k-1)$th time this has happened. In other words, $\eta_{1,i,1}$ is the number of stochastic gradient computations until the moment when the algorithm receives a “lucky” stochastic gradient where the last coordinate is not zeroed out. The random variable $\eta_{1,i,2}$ is the number of computations until it happens for the second time, and so on. Similarly, $\mu_{1,i,1}$ is the position of the first coordinate from the stream sent by the server to worker $i$ that belongs to $[K]$. The random variable $\mu_{1,i,2}$ refers to the second time this occurs, and so on. By definition, the sequences $\{\eta_{1,i,k}\}$ and $\{\mu_{1,i,k}\}$ follow approximately geometric-like distributions with parameters $p_{\sigma}$ and $p_{K},$ respectively.
@@
5 Lower Bound with Both W2S and S2W Communication
In the previous section, we provide the lower bound without taking into account the communication cost $\tau_{\textnormal{w}}.$ Combining Theorem 4.2 with our new Theorem D.1, which extends the result by Tyurin et al. (2024) for our setup, we can obtain the complete lower bound (5) from Theorem 1.6 with $\tau_{\textnormal{w}}>0$ and $\tau_{\textnormal{s}}>0.$ Notice that if $\tau_{\textnormal{s}}\simeq\tau_{\textnormal{w}},$ then the lower bound is $\tilde{\Omega}\left(\min\left\{h\left(\frac{\sigma^{2}}{n\varepsilon}+1\right)\frac{L\Delta}{\varepsilon}+{\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}\tau_{\textnormal{s}}d\frac{L\Delta}{\varepsilon}},h\frac{L\Delta}{\varepsilon}+{\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}h\frac{\sigma^{2}L\Delta}{\varepsilon^{2}}}\right\}\right).$ Up to logarithmic factors, under Assumptions 1.1, 1.2, 1.3, and 1.4, it is infeasible to improve both $d$ and $\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ as $n$ increases.
5.1 Algorithms almost matching the lower bound
6 Conclusion
We prove nearly tight lower bounds for centralized distributed optimization under the computation and communication Assumption 1.4. We show that even in the homogeneous scenario, it is not possible to scale both $d$ and $\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ by more than poly-logarithmic factors in $n.$ The lower bounds are tight only up to logarithmic factors. Thus, a possible challenging direction is to improve the powers of the logarithms, or even eliminate the logarithms entirely. The latter (if at all possible) may be very challenging and would likely require entirely different constructions and techniques. Another limitation is that the lower bounds are constructed using random sparsifiers. Due to the uncertainty principle (Safaryan et al., 2022), we conjecture that the bounds also hold for the entire family of unbiased compressors, but proving this would require more sophisticated constructions.
[…]
Appendix A Auxiliary Facts and Notations
[…]
@@ Appendix A Auxiliary Facts and Notations
A.1 Notations
:=$\mathbb{N}:=\{1,2,\dots\};$ $\left\|x\right\|$ is the output of the standard Euclidean norm for all $x\in\mathbb{R}^{d}$; $\left\langle x,y\right\rangle=\sum_{i=1}^{d}x_{i}y_{i}$ is the standard dot product; $\left\|A\right\|$ is the standard spectral/operator norm for all $A\in\mathbb{R}^{d\times d};$ $g=\mathcal{O}(f):$ exist $C>0$ such that $g(z)\leq C\times f(z)$ for all $z\in\mathcal{Z};$ $g=\Omega(f):$ exist $C>0$ such that $g(z)\geq C\times f(z)$ for all $z\in\mathcal{Z};$ $g=\Theta(f):$ $g=\mathcal{O}(f)$ and $g=\Omega(f);$ $g=\widetilde{\mathcal{O}}(f),g=\widetilde{\Omega}(f),g=\widetilde{\Theta}(f):$ the same as $g=\mathcal{O}(f),g=\Omega(f),g=\Theta(f),$ respectively, but up to logarithmic factors; $g\simeq h:$ $g$ and $h$ are equal up to universal positive constants; $g\gtrsim h:$ $g$ greater or equal to $h$ up to universal positive constants.
Appendix B Lower Bound
B.1 New Construction
@@
[…]
@@
B.2 Auxiliary Lemmas
In this section, we list useful properties of the functions $\Phi,$ $\Gamma,$ $\Psi_{a},$ and $F_{T,K,a}.$ We prove them in Section B.3.
Lemma B.1 (Carmon et al. (2020)).
@@ B.2 Auxiliary Lemmas
for all $x\in\mathbb{R}.$ Moreover, $\Phi^{\prime}(x)>1$ for all $-1<x<1.$
Lemma B.2.
@@
Lemma B.4.
For all $x\in\mathbb{R}^{T},$ $\textnormal{prog}^{1}(\nabla F_{T,K,a}(x))\leq\max\{\textnormal{prog}^{K}(x)+1,\textnormal{prog}^{1}(x)\}.$
Lemma B.5.
@@ B.3 Proof of lemmas
Since $p=\textnormal{prog}^{1}(x),$ the gradient of the last sum is zero because $\Gamma^{\prime}(0)=0.$ And the gradient of the third sum has a non-zero value in at most the $p$th coordinate. ∎
See B.5
@@
Since ${x_{j-1}}\geq 1,\dots,{x_{j-K}}\geq 1$ and $11>x_{j}>-$ 1 (see Lemmas B.1 and B.3), we get
@@
for all $i\neq j\in[T]$ due to Lemmas B.1 and B.3 and $e\geq a>1$
Notice that $\nabla^{2}F_{T,K,a}$ is $(2K+1)$–diagonal Hessian. Repeating a textbook analysis for completeness and denoting temporary ${\bm{H}}:=\nabla^{2}F_{T,K,a}$, we will show that
@@
where the second equality due to ${\bm{H}}$ is $(2K+1)$–diagonal. Using the Cauchy–Schwarz inequality,
@@
Appendix C Proof of Theorem 4.2
@@
(Step 1: Construction). Using the construction from Section B.1, we define a scaled version of it. Let us take any $\lambda>0,$ $d,T\in\mathbb{N},$ $d\geq T,$ and take the function $f\,:\,\mathbb{R}^{d}\to\mathbb{R}$ such that
@@
where $\ell_{1}(K,a)$ is defined in Lemma B.8 and $x_{[T]}\in\mathbb{R}^{T}$ is the vector with the first $T$ coordinates of $x\in\mathbb{R}^{d}.$ Notice that the last $d-T$ coordinates are artificial.
First, we have to show that $f$ is $L$-smooth and $f(0)-\inf_{x\in\mathbb{R}^{d}}f(x)\leq\Delta,$ Using Lemma B.8,
@@
where $\Delta^{0}(K,a)$ is defined in Lemma B.6. We also choose
@@
where the last inequality due to Lemma B.5. Note that
@@
$[\nabla f(x;\xi)]_{j}:=\nabla_{j}f(x)\left(1+\mathbbm{1}\left[j>\textnormal{prog}^{1}(x)\right]\left(\frac{\xi}{p_{\sigma}}-1\right)\right)\quad\forall x\in\mathbb{R}^{d},$ (26)
and $\mathcal{D}_{\xi}=\textnormal{Bernouilli}(p_{\sigma})$ for all $j\in[n],$ where $p_{\sigma}\in(0,1].$ We denote $[x]_{j}$ as the $j$th index of a vector $x\in\mathbb{R}^{d}.$ It is left to show this mapping is unbiased and $\sigma^{2}$-variance-bounded. Indeed,
${\mathbb{E}}\left[[\nabla f(x,\xi)]_{i}\right]=\nabla_{i}f(x)\left(1+\mathbbm{1}\left[i>\textnormal{prog}^{1}(x)\right]\left(\frac{{\mathbb{E}}\left[\xi\right]}{p_{\sigma}}-1\right)\right)=\nabla_{i}f(x)$
@@
where we use Lemma B.7 in the last inequality. Taking
@@
neither the server nor any worker is able to discover (filled with non-zero values) the $(K+1)$th and subsequent coordinates due to Lemma B.4.
@@
Option 1: In the first option, a worker computes a stochastic gradient, which takes $h$ seconds. However, due to the construction of stochastic gradients, even if the computation is completed, the worker will not make progress or discover a new non-zero coordinate, as it will be zeroed out with probability $p_{\sigma}$. Let $\eta_{1,i,1}$ be the number of stochastic gradients computations [^7] It is possible that $\mathbb{P}\left(\eta_{1,i,1}=\infty\right)>0$ if, for instance, the algorithm decides to stop calculating stochastic gradients. And even $\mathbb{P}\left(\eta_{1,i,1}=\infty\right)=1$ if it does not calculate at all. until the first moment when a coordinate is not zeroed out in (26) in worker $i.$ Assume that $\xi_{1},\xi_{2},\dots$ is a stream of i.i.d. random Bernoulli variables from (26) in worker $i$ (all workers have different streams), then
@@
for all $t\geq 0.$ Similarly, let $\eta_{1,i,k}$ denote the number of stochastic gradient computations until the first moment when a coordinate is not zeroed out in (26), after the moment when the $(k-1)$th coordinate is no longer zeroed out in worker $i.$ In other words, worker $i$ should calculate $\eta_{1,i,1}$ stochastic gradients to discover the first coordinate, calculate $\eta_{1,i,2}$ stochastic gradients to discover the second coordinate, and so on. Since the draws of $\xi$ in (26) are i.i.d., we can conclude that

### v2

@@ 2.3 Failure of the previous construction in the homogeneous setting
Indeed, if the server sends i.i.d. RandKK compressors with $K=1,$ then the number of consecutive coordinates that the server has to send before the workers receive the last non-zero coordinate is $\sum_{j=1}^{T}\min\limits_{i\in[n]}\eta_{ji},$ where $\mathbb{P}\left(\eta_{ji}=m\middle|\{\eta_{kj}\}_{k<i}\right)\leq p(1-p)^{m-1}$ for all $m,j\geq 1,i\in[n].$ The $\min\limits_{i\in[n]}$ operation appears because it is sufficient to wait for the first “luckiest” worker. Analyzing this sum, we can only show that
@@ 3 A New “Worst-Case” Function
In this section, we give a less technical description of our lower bound construction and the main theorem from Section D. Instead of (6), we propose to use another “worst-case” function. For any $T,K\in\mathbb{N},$ and $e\geq a>1,$ we define the function $F_{T,K,a}\,:\,\mathbb{R}^{T}\to\mathbb{R}$ such that
@@
$\Psi_{a}(x)=\begin{cases}0,&x\leq 1/2,\\ \exp\left(\log a\cdot\left(1-\frac{1}{(2x-1)^{2}}\right)\right),&x>1/2,\end{cases}\qquad\Phi(x)=\sqrt{e}\int_{-\infty}^{x}e^{-\frac{1}{2}t^{2}}dt,$ (10)
@@
and $x_{0}=\dots=x_{-K+1}\equiv 1.$ The main modification is that instead of the block $\Psi(x_{i-1})\Phi(x_{i}),$ we use $\Psi_{a}(x_{i-K})\dots\Psi_{a}(x_{i-2})\Psi_{a}(x_{i-1})\Phi(x_{i})$ (ignore $a$ for now). In the previous approach, it was sufficient for a worker to have $x_{i-1}\neq 0$ to discover the next $i$th coordinate. However, in our new construction, the worker needs $x_{i-1}\neq 0,x_{i-2}\neq 0,\dots,x_{i-K+1}\neq 0$ for that. With this modification, it is not sufficient for the “luckiest” worker to get the non-zero i1i-1th coordinate to discover the next coordinate: to progress to the $i$th coordinate, the worker should also have non-zero i2i-2th, …, iK+1i-K+1th coordinates.
@@
    For all $x\in\mathbb{R}^{T},$ $\textnormal{supp}(\nabla F_{T,K,a}(x))\subseteq\{1,\dots,\textnormal{prog}^{K}(x)+1\}\cup\textnormal{supp}(x),$
    where $\textnormal{supp}(v):=\{\,i\in[d]:v_{i}\neq 0\,\}.$
@@
Lemma 3.2 (Lemmas D.6, D.7, and D.8).
The function $F_{T,K,a}$ satisfies:
@@
Figure 1: The functions $\Psi_{a}(x)$ and $\Gamma(x)$, along with their derivatives $\Psi^{\prime}_{a}(x)$ and $\Gamma^{\prime}(x)$. The plots of $\Phi(x)$ and $\Phi^{\prime}(x)$ are shown in (Carmon et al., 2020).
@@ 4 Lower Bound with Server-to-Worker (S2W) Communication
Let $L,\Delta,\varepsilon,n,\sigma^{2},d,\tau_{\textnormal{s}},\tau_{\textnormal{w}},h>0$ be any numbers such that $\bar{c}_{1}\varepsilon\log^{4}(n+1)<L\Delta$ and dimension $d\geq\bar{c}_{3}\frac{L\Delta}{\log^{3}(n+1)\varepsilon}.$ Consider Protocol 1. For all $i\in[n]$ and $k\geq 0,$ compressor $\mathcal{C}^{k}_{i}$ selects and transmits $P^{k}_{i}$ uniformly random coordinates without replacement, scaled by any constants [^5] To potentially preserve unbiasedness. For instance, RandKK scales by $\nicefrac{{d}}{{K}}.$, where $P^{k}_{i}\in\{0,\dots,d\}$ may vary across each compressor  [^6] For instance, the compressors can be RandKK (see Def. C.1) with any $K\in[d]$, PermKK (Szlendak et al., 2021), Identity compressor when $P^{k}_{i}=d.$. Then, for any algorithm $A\in\mathcal{A}_{\textnormal{zr}}$ (Def. 2.3), there exists a function $f\,:\,\mathbb{R}^{d}\to\mathbb{R}$ such that $f$ is $L$-smooth, i.e., $\left\|\nabla f(x)-\nabla f(y)\right\|\leq L\left\|x-y\right\|$ for all $x,y\in\mathbb{R}^{d},$ and $f(0)-\inf_{x\in\mathbb{R}^{d}}f(x)\leq\Delta,$ exists a stochastic gradient oracles that satisfies Assumption 1.3, and ${\mathbb{E}}\left[\inf_{y\in S_{t}}\left\|\nabla f(y)\right\|^{2}\right]>\varepsilon$ for all
@@
The formulation of Theorem 4.2 is standard in the literature. However, following Tyurin and Richtárik (2023b), we present the lower bound in terms of time complexities rather than iteration complexities. Then, following Huang et al. (2022); He et al. (2023); Tyurin et al. (2024), we consider a subfamily of unbiased compressors based from Definition 1.5 on random sparsification to prove the lower bound; this is standard practice for taking the “worst-case” compressors from the family (similarly to taking the “worst-case” functions (Carmon et al., 2020; Nesterov, 2018)). Moreover, due to the uncertainty principle (Safaryan et al., 2022), all unbiased compressors exhibit variance and communication cost comparable to those of the RandKK sparsifier in the worst case (up to constant factors).
The main observation in Theorems 4.1 and 4.2 is that it is not possible to scale both $d$ and $\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ by more than $\log^{4}(n+1)$ and $\log^{6}(n+1),$ respectively. Asymptotically, this scaling is significantly worse than the linear $n$ and square-root $\sqrt{n}$ scalings discussed in Section 1.1. For instance, if n=10,000n=10{,}000 and $d$ is increased by a factor of $10$, we have to increase $n$ by a factor of $10^{3}$ (two factors more) to ensure that $\nicefrac{{\tau_{\textnormal{s}}d}}{{\log^{4}(n+1)}}$ does not change.
In Section A, we present the intuition and the proof sketch of the result.
5 Lower Bound with Both W2S and S2W Communication
In the previous section, we provide the lower bound without taking into account the communication cost $\tau_{\textnormal{w}}.$ Combining Theorem 4.2 with our new Theorem F.1, which extends the result by Tyurin et al. (2024) for our setup, we can obtain the complete lower bound (5) from Theorem 1.6 with $\tau_{\textnormal{w}}>0$ and $\tau_{\textnormal{s}}>0.$ Notice that if $\tau_{\textnormal{s}}\simeq\tau_{\textnormal{w}},$ then the lower bound is $\tilde{\Omega}\left(\min\left\{h\left(\frac{\sigma^{2}}{n\varepsilon}+1\right)\frac{L\Delta}{\varepsilon}+{\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}\tau_{\textnormal{s}}d\frac{L\Delta}{\varepsilon}},h\frac{L\Delta}{\varepsilon}+{\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}h\frac{\sigma^{2}L\Delta}{\varepsilon^{2}}}\right\}\right).$ Up to logarithmic factors, under Assumptions 1.1, 1.2, 1.3, and 1.4, it is infeasible to improve both $d$ and $\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ as $n$ increases.
5.1 Algorithms almost matching the lower bound
6 Conclusion
We prove nearly tight lower bounds for centralized distributed optimization under the computation and communication Assumption 1.4. We show that even in the homogeneous scenario, it is not possible to scale both $d$ and $\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ by more than poly-logarithmic factors in $n.$ Notice that the family of unbiased compressors contains the family of biased compressors (Beznosikov et al., 2020). Therefore, our lower bounds also apply to methods that use biased compressors, in the sense that there exists a “worst-case” compressor for which these methods cannot achieve a convergence rate faster than the lower bound in Theorem 1.6.
The lower bounds are tight only up to logarithmic factors. Thus, a possible challenging direction is to improve the powers of the logarithms, or even eliminate the logarithms entirely. The latter (if at all possible) may be very challenging and would likely require entirely different constructions and techniques. Another limitation is that the lower bounds are constructed using random sparsifiers. Due to the uncertainty principle (Safaryan et al., 2022), we conjecture that the bounds also hold for the entire family of unbiased compressors, but proving this would require more sophisticated constructions.
In practice, however, biased and unbiased compressors, including TopKK and RankKK (Alistarh et al., 2018; Vogels et al., 2019), exhibit significantly better compression properties than those predicted by worst-case analysis (Beznosikov et al., 2020). When used on the server side in combination with EF or EF21-P (Gruntkowska et al., 2023; Tyurin et al., 2024), they may help mitigate the pessimistic term $\color[rgb]{0.92578125,0.578125,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.578125,0.0546875}\tau_{\textnormal{s}}d\nicefrac{{L\Delta}}{{\varepsilon}}$}. Moreover, our pessimistic lower bound may potentially be broken under additional assumptions such as convexity or second-order smoothness.
Acknowledgements
The work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4F0002 and the agreement №139-10-2025-033.
Appendix A Proof Sketch
@@
Next, we define two sets of random variables: (i) let $\eta_{1,i,k}$ denote the number of stochastic gradient computations until the first moment when a coordinate is not zeroed out in the stochastic gradient oracle (see (26)), after the moment when the (k1)(k-1)th coordinate is no longer zeroed out in worker $i$; (ii) let $\mu_{1,i,k}$ be the number of received coordinates until the moment when the last received coordinate belongs to $[K]$, after the (k1)(k-1)th time this has happened. In other words, $\eta_{1,i,1}$ is the number of stochastic gradient computations until the moment when the algorithm receives a “lucky” stochastic gradient where the last coordinate is not zeroed out. The random variable $\eta_{1,i,2}$ is the number of computations until it happens for the second time, and so on. Similarly, $\mu_{1,i,1}$ is the position of the first coordinate from the stream sent by the server to worker $i$ that belongs to $[K]$. The random variable $\mu_{1,i,2}$ refers to the second time this occurs, and so on. By definition, the sequences $\{\eta_{1,i,k}\}$ and $\{\mu_{1,i,k}\}$ follow approximately geometric-like distributions with parameters $p_{\sigma}$ and $p_{K},$ respectively.
@@
Appendix B Additional Related Work
While we focus on lower bounds in the context of stochastic optimization and compressed vectors in nonconvex settings, there is much related work in other domains and setups. The seminal works on lower bounds were done by Nemirovskij and Yudin (1983); Nesterov (2018), where Nesterov (2018) showed that the accelerated gradient descent (Nesterov, 1983) is optimal in the convex setting using a quadratic “worst-case” function. In the nonconvex setting, Carmon et al. (2020) provided an alternative function, described in the main part of the paper. For convex problems, Woodworth et al. (2018) introduced the graph oracle, a generalization of the classical gradient oracle (Nemirovskij and Yudin, 1983; Nesterov, 2018), and established lower bounds for a broad class of parallel optimization methods. Arjevani et al. (2020b) further analyzed the delayed gradient descent method, which corresponds to Asynchronous SGD with constant iteration delays. Tyurin and Richtárik (2023b; 2024); Tyurin et al. (2024) proved lower bounds for methods in asynchronous settings. Fang et al. (2018); Patel et al. (2022) studied a different setting from Assumption 1.3, where they assumed the mean-squared smoothness property to enable the analysis of methods with variance reduction techniques (Fang et al., 2018; Cutkosky and Orabona, 2019). Woodworth and Srebro (2016) considered the finite-sum setting in the convex setting. Woodworth et al. (2020; 2021) proved that the min-max optimal algorithm for optimizing smooth convex objectives in the intermittent communication setting is the best of accelerated local and minibatch SGD, which leads to a similar conclusion to ours; however, their results are related to, but not directly comparable with ours, since we analyze the limited scalability of improving both stochastic noise and communication complexity through compressors. Glasgow et al. (2022) provided sharp lower bounds for local SGD approaches in terms of iteration complexity. Huang et al. (2022); He et al. (2023); Gruntkowska et al. (2024) provided lower bounds for compression techniques, but in the heterogeneous setting.
Appendix C Auxiliary Facts and Notations
Definition C.1 (RandKK).
Assume that $S$ is a random subset of $[d]$ such that $|S|=K$ for some $K\in[d]$. A stochastic mapping $\mathcal{C}\,:\,\mathbb{R}^{d}\times\mathbb{S}_{\nu}\rightarrow\mathbb{R}^{d}$ is called RandKK if
@@ Appendix C Auxiliary Facts and Notations
C.1 Notations
$\mathbb{N}:=\{1,2,\dots\};$ $\left\|x\right\|$ is the output of the standard Euclidean norm for all $x\in\mathbb{R}^{d}$; $\left\langle x,y\right\rangle=\sum_{i=1}^{d}x_{i}y_{i}$ is the standard dot product; $\left\|A\right\|$ is the standard spectral/operator norm for all $A\in\mathbb{R}^{d\times d};$ $g=\mathcal{O}(f):$ exist $C>0$ such that $g(z)\leq C\times f(z)$ for all $z\in\mathcal{Z};$ $g=\Omega(f):$ exist $C>0$ such that $g(z)\geq C\times f(z)$ for all $z\in\mathcal{Z};$ $g=\Theta(f):$ $g=\mathcal{O}(f)$ and $g=\Omega(f);$ $g=\widetilde{\mathcal{O}}(f),g=\widetilde{\Omega}(f),g=\widetilde{\Theta}(f):$ the same as $g=\mathcal{O}(f),g=\Omega(f),g=\Theta(f),$ respectively, but up to logarithmic factors; $\simeq h: gg$ and $h$ are equal up to universal positive constants; $\gtrsim h: gg$ greater or equal to $h$ up to universal positive constants; $\mathcal{C}$ is an unbiased compressor (Definition 1.5); $\textnormal{supp}(v)=\{\,i\in[d]:v_{i}\neq 0\,\};$ $h:$ maximum time (in seconds) for any worker to compute one stochastic gradient; $\tau_{\textnormal{s}}:$ communication time per coordinate from the server to any worker; $\tau_{\textnormal{w}}:$ communication time per coordinate from any worker to the server;
Appendix D Lower Bound
D.1 New Construction
@@
$\Psi_{a}(x)=\begin{cases}0,&x\leq 1/2,\\ \exp\left(\log a\cdot\left(1-\frac{1}{(2x-1)^{2}}\right)\right),&x>1/2,\end{cases}\qquad\Phi(x)=\sqrt{e}\int_{-\infty}^{x}e^{-\frac{1}{2}t^{2}}dt,$
@@
D.2 Auxiliary Lemmas
In this section, we list useful properties of the functions $\Phi,$ $\Gamma,$ $\Psi_{a},$ and $F_{T,K,a}.$ We prove them in Section D.3.
Lemma D.1 (Carmon et al. (2020)).
@@ D.2 Auxiliary Lemmas
for all $x\in\mathbb{R}.$ Moreover, $\Phi^{\prime}(x)>1$ for all 1<x<1.-1<x<1.
Lemma D.2.
@@
Lemma D.4.
For all $x\in\mathbb{R}^{T},$ $\textnormal{supp}(\nabla F_{T,K,a}(x))\subseteq\{1,\dots,\textnormal{prog}^{K}(x)+1\}\cup\textnormal{supp}(x),$ where $\textnormal{supp}(v):=\{\,i\in[d]:v_{i}\neq 0\,\}.$
Lemma D.5.
@@ D.3 Proof of lemmas
Since $p=\textnormal{prog}^{1}(x),$ the gradient of the last sum is zero because $\Gamma^{\prime}(0)=0.$ Moreover, if $x_{i}=0,$ then $\Gamma^{\prime}(x_{i})=0;$ thus, $\nabla\left(\sum_{i=1}^{p}\Gamma(x_{i})\right)\in\textnormal{supp}(x).$
See D.5
@@
Since ${x_{j-1}}\geq 1,\dots,{x_{j-K}}\geq 1$ and $1>x_{j}>-1$ (see Lemmas D.1 and D.3), we get
@@
for all $i\neq j\in[T]$ due to Lemmas D.1 and D.3 and $e\geq a>1$
Notice that $\nabla^{2}F_{T,K,a}$ is (2K+1)(2K+1)–diagonal Hessian. Repeating a textbook analysis for completeness and denoting temporary ${\bm{H}}:=\nabla^{2}F_{T,K,a}$, we will show that
@@
where the second equality due to $\bm{H}$} is (2K+1)(2K+1)–diagonal. Using the Cauchy–Schwarz inequality,
@@
Appendix E Proof of Theorem 4.2
@@
(Step 1: Construction). Using the construction from Section D.1, we define a scaled version of it. Let us take any $\lambda>0,$ $\in\mathbb{N}, dT,d\geq T$, and take the function $f\,:\,\mathbb{R}^{d}\to\mathbb{R}$ such that
@@
where $\ell_{1}(K,a)$ is defined in Lemma D.8 and $x_{[T]}\in\mathbb{R}^{T}$ is the vector with the first $T$ coordinates of $x\in\mathbb{R}^{d}.$ Notice that the last dTd-T coordinates are artificial.
First, we have to show that $f$ is $L$-smooth and $f(0)-\inf_{x\in\mathbb{R}^{d}}f(x)\leq\Delta,$ Using Lemma D.8,
@@
where $\Delta^{0}(K,a)$ is defined in Lemma D.6. We also choose
@@
where the last inequality due to Lemma D.5. Note that
@@
$[\nabla f(x;\xi)]_{j}:=\nabla_{j}f(x)\left(1+\mathbbm{1}\left[j=\textnormal{prog}^{K}(x)+1\right]\left(\frac{\xi}{p_{\sigma}}-1\right)\right)\quad\forall x\in\mathbb{R}^{d},$ (26)
and $\mathcal{D}_{\xi}=\textnormal{Bernouilli}(p_{\sigma}),$ where $p_{\sigma}\in(0,1].$ We denote $x]_{j}$ as the $j$th index of a vector $x\in\mathbb{R}^{d}.$ It is left to show this mapping is unbiased and $\sigma^{2}$-variance-bounded. Indeed,
${\mathbb{E}}\left[[\nabla f(x,\xi)]_{i}\right]=\nabla_{i}f(x)\left(1+\mathbbm{1}\left[i=\textnormal{prog}^{K}(x)+1\right]\left(\frac{{\mathbb{E}}\left[\xi\right]}{p_{\sigma}}-1\right)\right)=\nabla_{i}f(x)$
@@
where we use Lemma D.7 in the last inequality. Taking
@@
neither the server nor any worker is able to discover (filled with non-zero values) the $(K+1)$th and subsequent coordinates due to Lemma D.4.
@@
Option 1: In the first option, a worker computes a stochastic gradient, which takes $h$ seconds. However, due to the construction of stochastic gradients, even if the computation is completed, the worker will not make progress or discover a new non-zero coordinate, as it will be zeroed out with probability $p_{\sigma}$. Due to Lemma D.4, each worker can discover at most one coordinate at position $\textnormal{prog}^{K}(x)+1$ before time $t_{1}$ in the first $K$ coordinates, where $x$ is a query point.
Remark E.1.
For this reason, making multiple queries with the same random variable instead of a single query does not help the algorithm progress: if the coordinates are zeroed out, then they are zeroed out in all vectors.
Let $\eta_{1,i,1}$ be the number of stochastic gradients computations [^7] It is possible that $\mathbb{P}\left(\eta_{1,i,1}=\infty\right)>0$ if, for instance, the algorithm decides to stop calculating stochastic gradients. And even $\mathbb{P}\left(\eta_{1,i,1}=\infty\right)=1$ if it does not calculate at all. until the first moment when a coordinate is not zeroed out in (26) in worker $i.$ Assume that $\xi_{1},\xi_{2},\dots$ is a stream of i.i.d. random Bernoulli variables from (26) in worker $i$ (all workers have different streams), then
@@
for all $t\geq 0.$ Similarly, let $\eta_{1,i,k}$ denote the number of stochastic gradient computations until the first moment when a coordinate is not zeroed out in (26), after the moment when the (k1)(k-1)th coordinate is no longer zeroed out in worker $i.$ In other words, worker $i$ should calculate $\eta_{1,i,1}$ stochastic gradients to discover the first coordinate, calculate $\eta_{1,i,2}$ stochastic gradients to discover the second coordinate, and so on. Since the draws of $\xi$ in (26) are i.i.d., we can conclude that